calculate the number of molecules of oxalic acid in 100ml of 0.2 normality oxalic acid solution

Grade:12th pass

1 Answers

swetan
6 years ago
As oxalic acid is dibasic in nature,n-factor (bascity)=2Now,Normality=No .of moles/n-factor/volume in litres so,0.02=n×2÷0.1 I. e. n=0.001 thus no.of molecules=avagafro no/no.of moles= 6.022×10^20